COMMISSION 6 : ASTRONOMICAL TELEGRAMS (TELEGRAMMES ASTRONOMIQUES)

Report of Meeting, 24 July 1991

President: E. Roemer Secretary: B. G. Marsden

On opening the meeting, President Roemer requested participants to stand for a few moments in memory of the three members who had died since the previous meeting: Leland E. Cunningham, Edgar Everhart and Dmitrij Ya. Martynov. She noted that several non-members—but only four members—were present and that extra copies of the Commission’s report in IAU Trans. XXIA were available. Jinming Liu and D. J. Tholen were approved as new members of the Commission. J. E. Grindlay and R. M. West were confirmed as incoming President and Vice President for 1991-1994, with S. Isobe, B. G. Marsden and E. Roemer as the members of the Organizing Committee. D. W. E. Green, Associate Director of the Central Telegram Bureau, was made a Consultant.

Director Marsden reported that the Central Telegram Bureau had been functioning smoothly and that apart from the elimination of the old mechanical TWX machines in Sept. 1988 there had been no essential changes in operation since the Commission’s previous meeting. With the appearance of SN 1987A and the availability of the IAU Circulars by “setting host” over SPAN, followed in mid-1988 by the regular distribution by electronic mail, the rate of publication increased to what is now a rather steady 250 issues per year. However, subscriptions were converted from a per-issue to a per-month charge in 1986, when the rate at which Circulars were being issued was only about half what it is now. The printed version of the Circulars still commands the vast majority of the subscribers, and the bill for photo-offset printing has therefore more than doubled since the subscription rates were set; furthermore, although almost all the material received for the Circulars arrives in computer form, the preparation of them in a timely manner at all hours of the day and many of the night, during the week and at weekends, often involves considerable effort. A fairly substantial increase, 33\frac{1}{3} percent, in the subscription rate for the printed edition was therefore made in Sept. 1990. This increase was made reluctantly, for the deficit could have been eliminated if all those who see the electronic version of the Circulars were actually to take out subscriptions. It is unfortunate that extensive pirating occurs, with copies of the electronic versions of the Circulars being widely distributed to astronomers (principally professional) over much of the world, and the concomitant increase in the electronic subscription rates by 66\frac{2}{3} percent is an added burden only for the 200 or so honest electronic users. The number of electronic subscribers does continue to climb, some of the recent increase having come from amateur astronomers who receive the electronic Circulars at CompuServe addresses.

It is not easy to state a precise figure for the Bureau’s annual income, for the subscriptions are combined with those for the Minor Planet Circulars, and in any case, the electronic component is conducted as a joint venture with the Minor Planet Center. A working figure might be $100 000, some $20 000 of which is then spent on printing and another $20 000 of which concerns subscriptions to the Bureau’s telegram service, essentially at cost. A third $20 000 goes to the Smithsonian Institution in overhead charges, but this does not come close to meeting the Smithsonian’s contribution through postage, computer equipment, office rent and the salary of the Director. The remaining $40 000 goes toward salary and benefits for the Associate Director, as well as for secretarial and occasional student assistance and for supplies, some of which are also shared with the Minor Planet Center.

In the ensuing discussion, Isobe asked whether service to subscribers known to be making illicit redisseminations of the Circulars could be refused, and Tholen enquired about the copyright of the material. To copyright the material would not be practical and a disservice to astronomers, and summarily to cut off a subscriber who appeared to have paid in good faith should only be done as a very last resort. After all, the purpose of the Circulars is to inform all astronomers of new discoveries and other urgent information as rapidly as possible. That does not imply that users with means should not bear some responsibility for helping support the service, however, and there was a consensus that the best solution would be to explore the possibility of charging a considerably higher rate to those who make extensive redistributions. D. Steel wondered whether some kind of corporate sponsorship might be sought.

West enquired whether the printed Circulars, which are now sometimes held for as much as a week before being printed, and then take at least as long for delivery, could be discontinued. Marsden responded that there were still 500 people who subscribe only to the printed version and that the revenue from them is greater than that from the 200 who also receive the electronic version. W. Liller suggested that postcard reminders should be sent to subscribers when it was time to renew. Marsden responded that the special low subscription rate was established specifically to reduce this kind of administrative burden; invoices can certainly be sent to anyone who needs a reminder, but the regular subscription rate, which is 66\frac{2}{3} percent higher, is then charged. The regular rate can perhaps be viewed as a “library” rate, but there is no prohibition on librarians who wish to avail themselves of the lower rate.

The matter of line charges was brought up and whether an inability to pay them played a role in deciding whether items should be published. Stressing that the line charges were an important part of the Bureau’s income, Marsden responded that the number of rejections for this reason was almost negligible, and that if there were any, they were due, not to inability to pay, but to reluctance to pay. Items from amateurs are exempted from line charges, but great selectivity is then exercised over what is published—mainly magnitude estimates of comets and novae used as fillers.

Noting that the IAU Secretariat had been slow to forward the subventions to the Central Bureau in recent years, it was agreed to request the General Secretary automatically to include an appropriate subvention in the IAU budget and to negotiate a payment schedule with the Bureau’s Director.

The following resolution was then adopted:

Commission 6,

noting the indispensable character of the service rendered to the international astronomical community by the Central Bureau for Astronomical Telegrams by rapid communication of critical information,

calls attention to the importance of the token subvention as a demonstration of the support of the IAU for this crucial activity and strongly urges its continuation.

Report of Meeting, 30 July 1991

Although originally planned as a separate meeting on the services—particularly the electronic services—of the Central Bureau, this meeting became a joint one with Commission 5 on electronic communication generally. In the portion of particular interest to Commission 6, Marsden noted that there had been since 1988 a considerable increase in the number of countries to which the IAU Circulars are sent by e-mail, the only notable holdouts now being Bulgaria, China, Czechoslovakia and Romania. The movement toward the INTERNET system, in which each address ends in a two-letter country code (apart from the U.S.), was generally welcomed, but in several cases subscribers sent in addresses that were not operative, and in others addresses simply would not function with the mailing system (involving UUNET) being used for the e-mail distribution of the Circulars. A particular problem involving both INTERNET and BITNET addresses in Germany (.DE) was confirmed by others at the meeting.

Some subscribers, notably in South Africa, remarked that electronic delivery of the Circulars sometimes took as long as a week. It was agreed that the electronic mail systems used by astronomers were still far from perfect, G. Winkler remarking that one got what one paid for, and astronomers make no financial contribution to the upkeep of INTERNET. Marsden agreed that some of the delay was in computers used by the Central Bureau and that he hoped this situation could be improved during the course of the next few months.